The length of a rectangle is 3in more than twice its width. If the perimeter of the rectangle is 36in, find the width of the rectangle.
(a)7in
(b)6in
(c)4in
(d)5in

Answer by tutorcecilia(2152)

Use the formula for the perimeter of a rectangle:
P=2(length+width)
P=36
Length = 3in more than twice its width = 2 x Width +3 = 2w+3
Width=w
.
P=2(l+w) [plug-in the values and solve for the width(w)]
36=2[(2w+3)+w]
36/2=2/2(2x+3+w)
18=2w+3+w
18=3w+3
18-3=3w
15=3w
15/3=3w/3
5=w
.
checking:
Length = 2w+3=2(5)+3=13
Width=5
.
36=2(13+5)
36=2(18)
36=36 [checks out]